SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (Dec.20, 2005)

Attendees: Rainer Bartoldus, Kris Chaisanguanthum, Al Eisner, Fang Fang, Selina Li, Rahmat Rahmat, Dong Su.

Operations

  • PEP is continuing take data. No issues with trigger.

EMT

  • All prescale codes are checked in. Dan is waiting for the new release to be available.

DCT/DCZ

  • Dct/Dch calibration at the test stand is now working properly. Selina and Xuedong found out the fibers were plugged into the wrong TSF boards and screwed up the calibration.
  • Su Dong pointed out that the DCH spikes due to electronic noises are gone by raise the noise threshold. He also pointed out the dropped dch hits may be caused by left right ambiguity.

Interface Board

  • The three Interface boards at the test stand all works fine, except one board might have problem at the output to zpd.

Online

  • Fang has removed the dependency of online DAQ on teststand and calibration. She did dct/dch calibration with the new codes at the teststand and the results are fine. Fang will test at IR2 when there is a chance.

L3

  • Arik found a bug in the track fit. The bug especially affects low momentum tracks. Event rate is expected to increase after the bug fixed. Rainer and Arik are working on the fix.
  • Al suggested to have the trigger rate plots for physics events such as hadron. Currently the plots are available only for bhabha events. His concern is if we have enough statistics to do so.

Data Quality and Monitoring

  • Rahmat is performing the normal QA checks for Run4 reprocessing. Nothing unusual was seen. He will monitor data quality during the holiday period.
  • Rahmat is working on the strip charts.
  • The fix of the error bars in the raw efficiency plots are in the next release. Kris will find out if Prafulla's fix is also in.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.